### v2

# Arithmetic dynamics and Generalized Fermat’s conjecture

Arithmetic dynamics and Generalized Fermat’s conjecture
Atsushi MORIWAKI School of General Education, Chubu University, 1200 Matsumoto-cho, Kasugai-shi, Aichi 487-8501, JAPAN moriwaki@fsc.chubu.ac.jp
(Date: 14/Jan/2025)
1. Introduction
Let $K$ be an arithmetic function field (i.e., $K$ is finitely generated over $\mathbb{Q}$}) and $S=((\Omega,\mathcal{A},\nu),\{|\raisebox{1.72218pt}{.}|_{\omega}\}_{\omega\in\Omega})$ be a proper adelic structure of $K$ such that $S$ has Northcott’s property and $\nu(\mathcal{A})\not\subseteq\{0,1\}$ (see [3, Proposition 2.7.8] for the existence of such adelic structure). Let $X$ be a geometrically integral projective scheme over $K$ and $L$ be an ample line bundle on $X$. Fix a positive integer $N_{0}$. Let $F=\{f_{N}\}_{N=N_{0}}^{\infty}$ be a sequence of endomorphisms of $X$ with the following properties:
  1. (1)
    $f_{N}^{*}(L)\simeq L^{\otimes d_{N}}$ for some integer $d_{N}\geqslant 2$.
  2. (2)
    $\lim_{N\to\infty}\deg(f_{N})=\infty$.
  3. (3)
    $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}$ for all $N,N^{\prime}\geqslant N_{0}$.
The sequence $F$ is called a system of endomorphisms compatible with $L$. We say $F$ is multiplicative (resp. additive) if $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{NN^{\prime}}$ (resp. $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{N+N^{\prime}}$) for all $N,N^{\prime}\geqslant N_{0}$. For example, the cases of Example 3.2 and Example 3.3 are multiplicative, and the case of Example 3.4 is additive. We can assign the height function $h_{F}$ on $X(K)$ to $F$ such that
$h_{F}(x)\geqslant 0\ \text{and}\ h_{F}(f_{N}(x))=d_{N}h_{F}(x)\quad(\forall x\in X(K))$
(c.f., Proposition 3.1). Let $Y$ be an equidimensional subscheme of $X$ and $Y_{N}:=f_{N}^{-1}(Y)$ ($N\geqslant N_{0}$). We say that $Y_{N}$ has Fermat’s property over $K$ if
$Y_{N}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}$
(see Remark 3.5 for the reason why we use the terminology “Fermat’s property”). In this paper, we would like to propose the following conjecture and give its evidences.
Conjecture 1.1 (Generalized Fermat’s conjecture).
If there is $N_{1}\geqslant N_{0}$ such that $Y_{N}(K)$ is finite for any integer $N\geqslant N_{1}$, then does there exist $N_{2}\geqslant N_{0}$ such that $Y_{N}$ has Fermat’s property for any $N\geqslant N_{2}$?
Theorem 1.2.
  1. (1)
    If $Y(K)$ is finite, then there exists $N_{2}$ such that $Y_{N}$ has Fermat’s property for any $N\geqslant N_{2}$.
  2. (2)
    If $F$ is additive and there is $N_{1}\geqslant N_{0}$ such that $Y_{N_{1}}(K)$ is finite, then there exists $N_{2}\geqslant N_{0}$ such that $Y_{N}$ has Fermat’s property for any $N\geqslant N_{2}$. Namely, Conjecture 1.1 holds for an additive system of endomorphisms compatible with $L$.
  3. (3)
    If $F$ is multiplicative and there is $p_{0}$ such that $Y_{p}(K)$ is finite for any prime $p$ with $p\geqslant p_{0}$, then
    $\lim_{m\to\infty}\frac{\{N_{0}\leqslant N\leqslant m\mid\text{$F_{N}$ has Fermat's property over $K$}\}}{m}=1.$
2. Endomorphism compatible with an ample line bundle
Let $f:X\to X$ be an endomorphism over $K$. Let $L$ be an ample line bundle on $X$. We say $f$ is compatible with $L$ if $\alpha:f^{*}(L)\simeq L^{\otimes d(f)}$ for some positive integer $d(f)$.
Proposition 2.1.
  1. (1)
    $d(f)^{\dim X}=\deg(f)$.
From now on, we assume that $d(f)$ is greater than or equal to 22.
  1. (2)
    There exists the unique metric family $\varphi_{f,\alpha}=\{\varphi_{f,\alpha,\omega}\}_{\omega\in\Omega}$ of $L$, called the canonical global compactification of $L$, such that, for each $\omega\in\Omega$, the isomorphism $\alpha_{\omega}:f_{\omega}^{*}(L_{\omega})\simeq L_{\omega}^{\otimes d(f)}$ extends to the isometry $f_{\omega}^{*}(L_{\omega},\varphi_{f,\alpha,\omega})\simeq(L_{\omega},\varphi_{f,\alpha,\omega})^{\otimes d(f)}$. Moreover, $(L,\varphi_{f,\alpha})$ is a nef adelic line bundle. In particular, $h_{f}(x)\geqslant 0$ for any $x\in X(K)$.
  2. (3)
    The height function $h_{\varphi_{f,\alpha}}$ with respect to $\varphi_{f,\alpha}$ does not depend on the choice of the isomorphism $\alpha$, so that $h_{\varphi_{f,\alpha}}$ is denoted by $h_{f}$.
  3. (4)
    For any $x\in X(K)$, $h_{f}(f(x))=d(f)h_{f}(x)$.
  4. (5)
    Let $g$ be another endomorphism of $X$ such that $g$ is compatible with $L$ and $d(g)\geqslant 2$. If $f\circ g=g\circ f$, then $h_{f}=h_{g}$. Moreover, $f\circ g$ is compatible with $L$, and $d(f\circ g)=d(f)d(g)$.
  5. (6)
    If we set $a(h_{f}):=\inf\{h_{f}(x)\mid\text{$x\in X(K)$ and $h_{f}(x)>0$}\}$, then $a(h_{f})>0$.
Proof.
(1) Indeed,
$d(f)^{\dim X}(L^{\dim X})=\Big((L^{\otimes d(f)})^{\dim X}\Big)=(f^{*}(L)^{\dim X})=\deg(f)(L^{\dim X}),$
as required.
(2) For the existence of $\varphi_{f,\alpha}$, we refer to [1, Proposition 2.5.11 (4)]. The second assertion follows from [3, Proposition 9.3.7].
(3) Let $\beta$ be another isomorphism $\beta:f^{*}(L)\simeq L^{\otimes d(f)}$. Then we can find $c\in K^{\times}$ such that $\beta=c\alpha$. Thus, by [3, Proposition 9.3.3], for each $\omega\in\Omega$,
$|\raisebox{1.72218pt}{.}|_{\varphi_{f,\beta,\omega}}=|c|_{\omega}^{-1/(d(f)-1)}|\raisebox{1.72218pt}{.}|_{\varphi_{f,\alpha,\omega}}.$
Therefore we obtain (3).
(4) Indeed,
$h_{f}(f(x))$ $=h_{(L,\varphi_{f,\alpha})}(f(x))=h_{f^{*}(L,\varphi_{f,\alpha})}(x)=h_{(L,\varphi_{f,\alpha})^{\otimes d(f)}}(x)$
$=d(f)h_{(L,\varphi_{f,\alpha})}(x)=d(f)h_{f}(x).$
(5) Fix isomorphisms $\alpha:f^{*}(L)\simeq L^{\otimes d(f)}$ and $\beta:g^{*}(L)\simeq L^{\otimes d(g)}$. Then, by [3, Proposition 9.3.4], we can find $r\in K^{\times}$ such that $|\raisebox{1.72218pt}{.}|_{\varphi_{g,\beta,\omega}}=|r|_{\omega}^{-1/(d(f)-1)(d(g)-1)}|\raisebox{1.72218pt}{.}|_{\varphi_{f,\alpha,\omega}}$. Thus the first assertion follows. Moreover,
$(f\circ g)^{*}(L)=f^{*}(g^{*}(L))\simeq f^{*}(L^{\otimes d(g)})=f^{*}(L)^{\otimes d(g)}\simeq(L^{\otimes d(f)})^{\otimes d(g)}=L^{\otimes d(f)d(g)}.$
Thus the second assertion follows.
(6) By [1, Proposition 6.2.3], for any $C\in{\mathbb{R}}$, the set $\{x\in X(K)\mid h_{f}(x)\leqslant C\}$ is finite. Thus (6) follows. ∎
The following lemma is a key of this paper.
Lemma $2.$
Let $S$ be a finite subset of $X(K)$. Let $f:X\to X$ and $g:X\to X$ be endomorphisms of $X$ such that $f$ and $g$ are compatible with $L$, $d(f),d(g)\in{\mathbb{Z}}_{\geqslant 2}$, and $f\circ g=g\circ f$. If $d(g)a(h_{f})>\max\{h_{f}(s)\mid s\in S\}$ and $g(x)\in S$ for $x\in X(K)$, then $h_{f}(x)=0$.
Proof.
If we assume that $h_{f}(x)>0$, then, by (4) and (5) in Proposition 2.1,
$\max\{h_{f}(x)\mid x\in S\}\geqslant h_{f}(g(x))=h_{g}(g(x))=d(g)h_{g}(x)=d(g)h_{f}(x)\geqslant d(g)a(h_{f}),$
which is a contradiction. ∎
3. Generalized Fermat’s conjecture
Let $F=\{f_{N}\}_{N=N_{0}}^{\infty}$ be a system of endomorphisms compatible with $L$.
Proposition 3.1.
  1. (1)
    The height function $h_{f_{N}}$ ($N\geqslant N_{0})$ does not depend on NN, so they are denoted by $h_{F}$.
  2. (2)
    For $N\geqslant N_{0}$, $h_{F}(f_{N}(x))=d_{N}h_{F}(x)$.
Proof.
(1) follows from (5) in Proposition 2.1.
(2) By (4) in Proposition 2.1,
$h_{F}(f_{N}(x))=h_{f_{N}}(f_{N}(x))=d_{N}h_{f_{N}}(x)=d_{N}h_{F}(x),$
as required. ∎
The following are examples of systems of endomorphisms compatible with $L$.
Example 3.2.
Let $X={\mathbb{P}}^{n}_{K}$, $L={\mathcal{O}}_{{\mathbb{P}}^{n}}(1)$ and $f_{N}(x_{0}:\cdots:x_{n})=(x_{0}^{N}:\cdots:x_{n}^{N})$ ($N\geqslant 2$). Then $f_{N}^{*}(L)\simeq L^{\otimes N}$ and $f_{N}\circ f_{N^{\prime}}=f_{N}\circ f_{N^{\prime}}=f_{NN^{\prime}}$. Note that $d_{N}=N$ and $\{f_{N}\}_{N=2}^{\infty}$ is multiplicative. In this case, $h_{F}(x_{0}:\cdots:x_{n})=0$ if and only if there is $t\in K^{\times}$ such that $tx_{0},\ldots,tx_{n}\in\{0\}\cup\mu(K)$, where $\mu(K)$ is the set of all roots of the unity in $K$.
Example $3.$
Let $X$ be an abelian variety over $K$, $L$ be an even ample line bundle on $X$, and $f_{N}(x)=Nx$ ($N\geqslant 2$). Then $f_{N}^{*}(L)\simeq L^{\otimes N^{2}}$ and $f_{N}\circ f_{N^{\prime}}=f_{N}\circ f_{N^{\prime}}=f_{NN^{\prime}}$. Note that $d_{N}=N^{2}$ and $\{f_{N}\}_{N=2}^{\infty}$ is multiplicative. In this case, $h_{F}(x)=0$ if and only if $x$ is a torsion point.
Example 3.4.
Let $f$ be an endomorphism of $X$ such that $f$ is compatible with $L$ and $d(f)\geqslant 2$. We set $f_{N}=f^{N}$ ($N\geqslant 1$). Then $(f_{N})^{*}(L)\simeq L^{\otimes d(f)^{N}}$ and $f_{N}\circ f_{N^{\prime}}=f_{N}\circ f_{N^{\prime}}=f_{N+N^{\prime}}$. Note that $d_{N}=d(f)^{N}$ and $\{f_{N}\}_{N=1}^{\infty}$ is additive. In this case, $h_{F}(x)=0$ if and only if $x$ is preperiodic with respect to $f$.
Remark 3.5.
In Example 3.2, we assume that $K={\mathbb{Q}}$, $n=2$ and $Y$ is defined by $X_{0}+X_{1}-X_{2}=0$. Note that $Y_{N}$ is given by $X_{0}^{N}+X_{1}^{N}-X_{2}^{N}$. If $Y_{N}$ has Fermat’s property, then, for $(x_{0}:x_{1}:x_{n})\in Y_{N}(K)$, we can find $t\in{\mathbb{Q}}^{\times}$ such that $tx_{0},tx_{1},tx_{2}\in\{0,1,-1\}$. Thus we can easily see that $x_{0}x_{1}x_{2}=0$, and hence Fermat’s conjecture for NN follows.
Remark 3.6.
If $\dim Y=1$, then $\dim Y_{N}=1$ because $f_{N}$ is finite. Thus if there is $N_{0}$ such that every irreducible component of $Y_{N}$ has the genus $\geqslant 2$ for any integer $N\geqslant N_{0}$, then, by Faltings theorem (c.f. [4]), $Y_{N}(K)$ is finite for $N\geqslant N_{0}$.
Proof of Theorem 1.2.
Let $S$ be a finite subset of $X(K)$. As $\lim_{m\to\infty}d_{m}=\infty$, there is $m_{0}$ such that
$d_{m}a(h_{F})>\max\{h_{F}(y)\mid y\in S\}$
for all $m\in{\mathbb{Z}}_{\geqslant m_{0}}$.
(1) We set $S=Y(K)$. Then, for $N\in{\mathbb{Z}}_{\geqslant m_{0}}$, as $f_{N}(Y_{N}(K))\subseteq S$ and $f_{N}\circ f_{N_{0}}=f_{N_{0}}\circ f_{N}$, by Lemma 2.2, $Y_{N}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}$.
(2) Note that $f^{-1}_{N}(Y_{N_{1}})=Y_{N+N_{1}}$. Thus the assertion follows from (1).
(3) For any prime number $p\geqslant p_{0}$, we set $S=Y_{p}(K)$. If $x\in Y_{mp}(K)$, then one has $f_{m}(x)\in Y_{p}(K)$ because $f_{p}(f_{m}(x))=f_{mp}(x)\in Y(K)$. Moreover, $f_{m}\circ f_{N_{0}}=f_{N_{0}}\circ f_{m}$. Thus, by Lemma 2.2, $Y_{mp}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}$ for $m\in{\mathbb{Z}}_{\geqslant m_{0}}$. Therefore, by [8, Lemma 5.16], we have the assertion. ∎
Remark 3.7.
Theorem 1.2 was previously discussed in the following cases:
  1. (1)
    $K={\mathbb{Q}}$, $X={\mathbb{P}}^{2}$, $f_{N}(x_{0}:x_{1}:x_{2})=(x_{0}^{N}:x_{1}^{N}:x_{2}^{N})$ and $Y$ is given by $X_{0}+X_{1}-X_{2}$ (M. Filaseta [5], A. Granville [6], D. R. Heath-Brown [7]).
  2. (2)
    $K$ is a number field, $X={\mathbb{P}}^{2}$, $f_{N}(x_{0}:x_{1}:x_{2})=(x_{0}^{N}:x_{1}^{N}:x_{2}^{N})$ and $Y$ is given by $X_{0}+X_{1}-X_{2}$ (H. Ikoma, S. Kawaguchi and A. Moriwaki [8]).
  3. (3)
    $K$ is an arithmetic function field, $X={\mathbb{P}}^{2}$, $f_{N}(x_{0}:x_{1}:x_{2})=(x_{0}^{N}:x_{1}^{N}:x_{2}^{N})$ and $Y$ is a strongly non-degenerated curve, that is, $Y_{N}$ is smooth over $K$ for all NN (H. Chen and A. Moriwaki [2]).
Example 3.8.
Let $X:={\mathbb{P}}^{1}\times{\mathbb{P}}^{1}(=\operatorname{Proj}(K[X_{0},X_{1}])\times\operatorname{Proj}(K[Y_{0},Y_{1}]))$, and $f_{N}:X\to X$ $(N\geqslant 2$) is defined to be
$f_{N}((x_{0}:x_{1})\times(y_{0}:y_{1}))=(x_{0}^{N}:x_{1}^{N})\times(y_{0}^{N}:y_{1}^{N}).$
Let $p_{i}:X\to{\mathbb{P}}^{1}$ be the projection to the $i$-th factor, and $L=p_{1}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))\otimes p_{2}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))$. Note that $f_{N^{\prime}}\circ f_{N}=f_{N}\circ f_{N^{\prime}}=f_{NN^{\prime}}$ (multiplicative), and that $f_{N}:X\to X$ is compatible with respect to $L$, that is, $f_{N}^{*}(L)=L^{\otimes N}$. Let $Y$ be a subvariety given by
$aX_{0}Y_{0}+bX_{1}Y_{1}+cX_{0}Y_{1}+dX_{1}Y_{0}=(X_{0}\ Y_{0})A\begin{pmatrix}Y_{0}\\ Y_{1}\end{pmatrix}=0\quad(a,b,c,d\in K),$
where $A=\begin{pmatrix}a&c\\ d&b\end{pmatrix}$. Then $Y_{N}$ is given by
$aX_{0}^{N}Y_{0}^{N}+bX_{1}^{N}Y_{1}^{N}+cX_{0}^{N}Y_{1}^{N}+dX_{1}^{N}Y_{0}^{N}=(X^{N}_{0}\ Y^{N}_{0})A\begin{pmatrix}Y^{N}_{0}\\ Y^{N}_{1}\end{pmatrix}=0.$
We assume that $abcd\not=0$ and $\det A\not=0$. Then it is easy to see that $Y_{N}$ is smooth over $K$ for all NN, so its genus is $(N-1)^{2}$. Thus, (3) in Theorem 1.2 holds in this case.
Fix a positive integer $p\geqslant 2$. We set $g_{N}=f_{p^{N}}$ for $N\geqslant 1$. Then
$g_{N^{\prime}}\circ g_{N}=g_{N}\circ g_{N^{\prime}}=f_{p^{N}p^{N^{\prime}}}=f_{p^{N+N^{\prime}}}=g_{N+N^{\prime}},$
so $\{g_{N}\}_{N=1}^{\infty}$ is additive. Thus generalized Fermat’s conjecture holds for $\{g_{N}\}_{N=1}^{\infty}$.
Example 3.9.
Here we consider a family of Lattès maps. Let $E$ be an elliptic curve over $K$ and $\pi:E\to{\mathbb{P}}^{1}$ be the morphism such that $\deg(\pi)=2$ and $\pi(P)=\pi(-P)$, that is, ${\mathbb{P}}^{1}=E/\{\pm 1\}$. For an integer $N\geqslant 2$, let $\varphi_{N}:{\mathbb{P}}^{1}\to{\mathbb{P}}^{1}$ be the morphism such that the following diagram is commutative:
[…]
Then $\varphi_{N}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))={\mathcal{O}}_{{\mathbb{P}}^{1}}(N^{2})$ and $\varphi_{N}\circ\varphi_{N^{\prime}}=\varphi_{N^{\prime}}\circ\varphi_{N}=\varphi_{NN^{\prime}}$. Thus if we define $F_{N}:{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}\to{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ to be $f_{N}=\varphi_{N}\times\varphi_{N}$, then we have a system of endomorphisms compatible with $L=p_{1}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))\otimes p_{2}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))$, where $p_{i}:X\to{\mathbb{P}}^{1}$ is the projection to the $i$-th factor, so the same assertions hold as in Example 3.8.