### v4

# Arithmetic dynamics and Generalized Fermat’s conjecture

Abstract.
In this short paper, we would like to propose generalized Fermat’s conjecture over an arithmetic function field in the framework of arithmetic dynamics. We also give several evidences for the conjecture.
I would like to thank Prof. Kawaguchi for his helpful comments. I also appreciate the anonymous useful suggestions.
1. Introduction
Let $K$ be an arithmetic function field (i.e., $K$ is finitely generated over $\mathbb{Q}$}) and $S=((\Omega,\mathcal{A},\nu),\{|\raisebox{1.72218pt}{.}|_{\omega}\}_{\omega\in\Omega})$ be a proper adelic structure of $K$ with Northcott’s property and $\nu(\mathcal{A})\not\subseteq\{0,1\}$ (for the definition of “adelic structure” and its related terminology, see Section 2). Let $X$ be a geometrically integral projective scheme over $K$ and $L$ be an ample line bundle on $X$. Fix a positive integer $N_{0}$. Let $F=\{f_{N}\}_{N=N_{0}}^{\infty}$ be a sequence of endomorphisms of $X$ with the following properties:
  1. (1)
    $f_{N}^{*}(L)\simeq L^{\otimes d_{N}}$ for some integer $d_{N}\geqslant 2$.
  2. (2)
    $\lim_{N\to\infty}\deg(f_{N})=\infty$.
  3. (3)
    $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}$ for all $N,N^{\prime}\geqslant N_{0}$.
The sequence $F$ is called a system of endomorphisms polarized by $L$. We say $F$ is multiplicative (resp. additive) if $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{NN^{\prime}}$ (resp. $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{N+N^{\prime}}$) for all $N,N^{\prime}\geqslant N_{0}$. For example, the cases of Example 4.2, Example 4.3, Example 4.4 and Example 4.5 are multiplicative, and the case of Example 4.6 is additive. We can assign the height function $h_{F}$ on $X(K)$ to $F$ such that
$h_{F}(x)\geqslant 0\ \text{and}\ h_{F}(f_{N}(x))=d_{N}h_{F}(x)\quad(\forall x\in X(K))$
(c.f., Proposition 4.1). Let $Y$ be an equidimensional subscheme of $X$ and $Y_{N}:=f_{N}^{-1}(Y)$ ($N\geqslant N_{0}$). We say that $Y_{N}$ has Fermat’s property over $K$ if
$Y_{N}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}$
(see Remark 4.8 for the reason why we use the terminology “Fermat’s property”). In this paper, we would like to propose the following conjecture and give its evidences.
Conjecture 1.1 (Generalized Fermat’s conjecture).
If there is $N_{1}\geqslant N_{0}$ such that $Y_{N}(K)$ is finite for any integer $N\geqslant N_{1}$, then does there exist $N_{2}\geqslant N_{0}$ such that $Y_{N}$ has Fermat’s property over $K$ for any $N\geqslant N_{2}$? For a multi-indexed version, see Conjecture 5.4.
Theorem 1.2.
  1. (1)
    If $Y(K)$ is finite, then there exists $N_{2}$ such that $Y_{N}$ has Fermat’s property for any $N\geqslant N_{2}$.
  2. (2)
    If $F$ is additive and there is $N_{1}\geqslant N_{0}$ such that $Y_{N_{1}}(K)$ is finite, then there exists $N_{2}\geqslant N_{0}$ such that $Y_{N}$ has Fermat’s property for any $N\geqslant N_{2}$. Namely, Conjecture 1.1 holds for an additive system of endomorphisms polarized by $L$.
  3. (3)
    If $F$ is multiplicative and there is a positive integer $p_{0}$ such that $Y_{p}(K)$ is finite for any prime $p$ with $p\geqslant p_{0}$, then
    $\lim_{m\to\infty}\frac{\#\{N_{0}\leqslant N\leqslant m\mid\text{$F_{N}$ has Fermat's property over $K$}\}}{m}=1,$
    that is, Fermat’s property holds with probability one.
2. Adelic structure of field
In this section, we quickly review several notions related to an adelic structure of a field. For details, see the manuscripts [3, 4, 5] due to H. Chen and A. Moriwaki.
An adelic structure of a field $K$ consists of data $S=((\Omega,\mathcal{A},\nu),\{|\raisebox{1.72218pt}{.}|_{\omega}\}_{\omega\in\Omega})$ satisfying the following properties:
  1. (1)
    $(\Omega,\mathcal{A},\nu)$ is a measure space, that is, $\mathcal{A}$ is a $\sigma$-algebra of $\Omega$ and $\nu$ is a measure on the measurable space $(\Omega,\mathcal{A})$.
  2. (2)
    $\{|\raisebox{1.72218pt}{.}|_{\omega}\}_{\omega\in\Omega}$ is a collection of absolute values of $K$ indexed by $\Omega$.
  3. (3)
    For $a\in K^{\times}$, the function $(\omega\in\Omega)\mapsto\log|a|_{\omega}$ is $\nu$-integrable.
For $\omega\in\Omega$, the completion of $K$ with respect to $|\raisebox{1.72218pt}{.}|_{\omega}$ is denoted by $K_{\omega}$. The adelic structure $S$ is said to be proper if
(2.1) $\int_{\Omega}\log|a|_{\omega}\nu(\mathrm{d}\omega)=0$
holds for all $a\in K^{\times}$. The equation (2.1) is called the product formula. From now on, we assume that $S$ is proper.
[…]
$\int_{\Omega}\log\|s\|_{\omega}\nu(d\omega)$
[…]
Let $X$ be a geometric integral projective scheme over $K$, $X_{\omega}:=X\times_{\operatorname{Spec}K}\operatorname{Spec}K_{\omega}$ for $\omega\in\Omega$, and $X_{\omega}^{\mathrm{an}}$ be the analytification in the sense of Berkovich [2]. Let $L$ be an invertible sheaf on $X$ and $L_{\omega}^{\mathrm{an}}$ be the pull-back of $L$ via the natural map $X_{\omega}^{\mathrm{an}}\to X$. For each $\omega\in\Omega$, let $\varphi_{\omega}$ be a continuous metric of $L_{\omega}^{\mathrm{an}}$. The collection $\varphi=\{\varphi_{\omega}\}_{\omega\in\Omega}$ is said to be adelic if it satisfies the conditions described in [3, Chapter 6]. The exact definition of the adelic collection is complicated in a few words, but note that, if $\varphi$ is adelic, then, for $x\in X(K)$ (i.e., a morphism $x:\operatorname{Spec}(K)\to X$ over $K$), the pullback $x^{*}(L,\varphi)$ yields a one-dimensional vector space over $K$ with an adelic norm family. Thus the height of $x$ with respect to $(L,\varphi)$ is defined by $\widehat{\operatorname{deg}}(x^{*}(L,\varphi))$ which is denoted by $h_{(L,\varphi)}(x)$.
We say $S$ has Northcott’s property if, for any $C\in{\mathbb{R}}$, the set
$\left\{a\in K\ \left|\ \int_{\Omega}\log\max\{|a|_{\omega},1\}\nu(d\omega)\leqslant C\right\}\right$.
is finite. Note that, by [3, Proposition 6.2.3], if $S$ has Northcott’s property and $L$ is ample, then the set $\{x\in X(K)\mid h_{(L,\varphi)}(x)\leqslant C\}$ is finite for all $C\in{\mathbb{R}}$. Moreover, if $K$ is an arithmetic function field, then, by [4, Theorem 2.7.18], $K$ has a proper adelic structure with Northcott’s property and $\nu(\mathcal{A})\not\subseteq\{0,1\}$. Finally note that once we have a proper adelic structure with Northcott’s property, the theory of height functions works similarly to the case of number fields.
3. Endomorphism polarized by an ample line bundle
Let $f:X\to X$ be an endomorphism over $K$. Let $L$ be an ample line bundle on $X$. We say $f$ is polarized by $L$ if there is an isomorphism $\alpha:f^{*}(L)\simeq L^{\otimes d(f)}$ for some positive integer $d(f)$.
Proposition 3.1.
  1. (1)
    $d(f)^{\dim X}=\deg(f)$.
From now on, we assume that $d(f)$ is greater than or equal to $2$.
  1. (2)
    There exists the unique metric family $\varphi_{f,\alpha}=\{\varphi_{f,\alpha,\omega}\}_{\omega\in\Omega}$ of $L$, called the canonical global compactification of $L$, such that, for each $\omega\in\Omega$, the isomorphism $\alpha_{\omega}:f_{\omega}^{*}(L_{\omega})\simeq L_{\omega}^{\otimes d(f)}$ extends to the isometry $f_{\omega}^{*}(L_{\omega},\varphi_{f,\alpha,\omega})\simeq(L_{\omega},\varphi_{f,\alpha,\omega})^{\otimes d(f)}$. Moreover, $(L,\varphi_{f,\alpha})$ is a nef adelic line bundle. In particular, $h_{f}(x)\geqslant 0$ for any $x\in X(K)$.
  2. (3)
    The height function $h_{\varphi_{f,\alpha}}$ with respect to $\varphi_{f,\alpha}$ does not depend on the choice of the isomorphism $\alpha$, so that $h_{\varphi_{f,\alpha}}$ is denoted by $h_{f}$.
  3. (4)
    For any $x\in X(K)$, $h_{f}(f(x))=d(f)h_{f}(x)$.
  4. (5)
    Let $g$ be another endomorphism of $X$ such that $g$ is polarized by $L$ and $d(g)\geqslant 2$. If $f\circ g=g\circ f$, then $h_{f}=h_{g}$. Moreover, $f\circ g$ is polarized by $L$, and $d(f\circ g)=d(f)d(g)$.
  5. (6)
    If we set $a(h_{f}):=\inf\{h_{f}(x)\mid\text{$x\in X(K)$ and $h_{f}(x)>0$}\}$, then $a(h_{f})>0$.
Proof.
(1) Indeed,
$d(f)^{\dim X}(L^{\dim X})=\Big((L^{\otimes d(f)})^{\dim X}\Big)=(f^{*}(L)^{\dim X})=\deg(f)(L^{\dim X}),$
as required.
(2) For the existence of $\varphi_{f,\alpha}$, we refer to [3, Proposition 2.5.11 (4)]. The second assertion follows from [5, Proposition 9.3.7].
(3) Let $\beta$ be another isomorphism $\beta:f^{*}(L)\simeq L^{\otimes d(f)}$. Then we can find $c\in K^{\times}$ such that $\beta=c\alpha$. Thus, by [5, Proposition 9.3.3], for each $\omega\in\Omega$,
$|\raisebox{1.72218pt}{.}|_{\varphi_{f,\beta,\omega}}=|c|_{\omega}^{-1/(d(f)-1)}|\raisebox{1.72218pt}{.}|_{\varphi_{f,\alpha,\omega}}.$
Therefore we obtain (3).
(4) Indeed,
$h_{f}(f(x))$ $=h_{(L,\varphi_{f,\alpha})}(f(x))=h_{f^{*}(L,\varphi_{f,\alpha})}(x)=h_{(L,\varphi_{f,\alpha})^{\otimes d(f)}}(x)$
$=d(f)h_{(L,\varphi_{f,\alpha})}(x)=d(f)h_{f}(x).$
(5) Fix isomorphisms $\alpha:f^{*}(L)\simeq L^{\otimes d(f)}$ and $\beta:g^{*}(L)\simeq L^{\otimes d(g)}$. Let us consider the following homomorphisms:
$\begin{cases}\begin{CD}g^{*}(f^{*}(L))@>{\sim}>{{g^{*}(\alpha)}}>g^{*}(L^{\otimes d(f)})@>{\sim}>{\beta^{\otimes d(f)}}>L^{\otimes d(f)d(g)}\end{CD},\\[8.61108pt] \begin{CD}g^{*}(f^{*}(L))=f^{*}(g^{*}(L))@>{\sim}>{f^{*}(\beta)}>f^{*}(L^{\otimes d(g)})@>{\sim}>{\alpha^{\otimes d(g)}}>L^{\otimes d(f)d(g)}\end{CD}.\end{cases}$
Thus we can find $r\in K^{\times}$ such that $\beta^{\otimes d(f)}\circ g^{*}(\alpha)=r\cdot\alpha^{\otimes d(g)}\circ f^{*}(\beta)$. Then, by [5, Proposition 9.3.4], one has $|\raisebox{1.72218pt}{.}|_{\varphi_{g,\beta,\omega}}=|r|_{\omega}^{-1/(d(f)-1)(d(g)-1)}|\raisebox{1.72218pt}{.}|_{\varphi_{f,\alpha,\omega}}$. Thus the first assertion follows. Moreover, by using the above isomorphisms, the second assertion is obtained.
(6) Note that, for any $C\in{\mathbb{R}}$, the set $\{x\in X(K)\mid h_{f}(x)\leqslant C\}$ is finite. Thus (6) follows. ∎
The following lemma is a key of this paper.
Lemma 3.2.
Let $S$ be a finite subset of $X(K)$. Let $f:X\to X$ and $g:X\to X$ be endomorphisms of $X$ such that $f$ and $g$ are polarized by $L$, $d(f),d(g)\in{\mathbb{Z}}_{\geqslant 2}$, and $f\circ g=g\circ f$. If $d(g)a(h_{f})>\max\{h_{f}(s)\mid s\in S\}$ and $g(x)\in S$ for $x\in X(K)$, then $h_{f}(x)=0$.
Proof.
If we assume that $h_{f}(x)>0$, then, by (4) and (5) in Proposition 3.1,
$\max\{h_{f}(x)\mid x\in S\}\geqslant h_{f}(g(x))=h_{g}(g(x))=d(g)h_{g}(x)=d(g)h_{f}(x)\geqslant d(g)a(h_{f}),$
which is a contradiction. ∎
4. Generalized Fermat’s conjecture
Let $F=\{f_{N}\}_{N=N_{0}}^{\infty}$ be a system of endomorphisms polarized by $L$.
Proposition 4.1.
  1. (1)
    The height function $h_{f_{N}}$ ($N\geqslant N_{0})$ does not depend on $N$, so they are denoted by $h_{F}$.
  2. (2)
    For $N\geqslant N_{0}$, $h_{F}(f_{N}(x))=d_{N}h_{F}(x)$.
Proof.
(1) follows from (5) in Proposition 3.1.
(2) By (4) in Proposition 3.1,
$h_{F}(f_{N}(x))=h_{f_{N}}(f_{N}(x))=d_{N}h_{f_{N}}(x)=d_{N}h_{F}(x),$
as required. ∎
The following are examples of systems of endomorphisms polarized by $L$.
Example 4.2.
Let $X={\mathbb{P}}^{n}_{K}$, $L={\mathcal{O}}_{{\mathbb{P}}^{n}}(1)$ and $f_{N}(x_{0}:\cdots:x_{n})=(x_{0}^{N}:\cdots:x_{n}^{N})$ ($N\geqslant 2$). Then $f_{N}^{*}(L)\simeq L^{\otimes N}$ and $f_{N}\circ f_{N^{\prime}}=f_{N}\circ f_{N^{\prime}}=f_{NN^{\prime}}$. Note that $d_{N}=N$ and $\{f_{N}\}_{N=2}^{\infty}$ is multiplicative. In this case, $h_{F}(x_{0}:\cdots:x_{n})=0$ if and only if there is $t\in K^{\times}$ such that $tx_{0},\ldots,tx_{n}\in\{0\}\cup\mu(K)$, where $\mu(K)$ is the set of all roots of the unity in $K$.
Example 4.3.
Let $X$ be an abelian variety over $K$, $L$ be an even ample line bundle on $X$, and $f_{N}(x)=Nx$ ($N\geqslant 2$). Then $f_{N}^{*}(L)\simeq L^{\otimes N^{2}}$ and $f_{N}\circ f_{N^{\prime}}=f_{N}\circ f_{N^{\prime}}=f_{NN^{\prime}}$. Note that $d_{N}=N^{2}$ and $\{f_{N}\}_{N=2}^{\infty}$ is multiplicative. In this case, $h_{F}(x)=0$ if and only if $x$ is a torsion point.
Example 4.4.
Here we consider a family of Lattès maps. Let $E$ be an elliptic curve over $K$ and $\pi:E\to{\mathbb{P}}^{1}$ be the morphism such that $\deg(\pi)=2$ and $\pi(P)=\pi(-P)$, that is, ${\mathbb{P}}^{1}=E/\{\pm 1\}$. For an integer $N\geqslant 2$, let $f_{N}:{\mathbb{P}}^{1}\to{\mathbb{P}}^{1}$ be the morphism such that the following diagram is commutative:
[…]
Then $f_{N}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))={\mathcal{O}}_{{\mathbb{P}}^{1}}(N^{2})$ and $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{NN^{\prime}}$. Note that $d_{N}=N^{2}$ and $\{f_{N}\}_{N=2}^{\infty}$ is multiplicative.
Example 4.5.
Let $T_{N}(z)$ be the Chebyshev polynomial of degee $N$, which is given by the recursive relation $T_{N}(z)=zT_{N-1}(z)-T_{N-2}(z)$ ($T_{0}=2$, $T_{1}=z$ and $T_{2}=z^{2}-2$). It can be characterized by $T_{N}(z+z^{-1})=z^{N}+z^{-N}$, so we can see that $T_{N}(T_{N^{\prime}}(z))=T_{N^{\prime}}(T_{N}(z))=T_{NN^{\prime}}(z)$. Thus $\{T_{N}\}_{N=2}^{\infty}$ yields a multiplicative system of endomorphisms. Note that $(1/2)T_{N}(2z)$ is the classical Chebyshev polynomial of degree $N$. For details, see [11, Section 6.2].
Example 4.6.
Let $f$ be an endomorphism of $X$ such that $f$ is polarized by $L$ and $d(f)\geqslant 2$. We set $f_{N}=f^{N}$ ($N\geqslant 1$). Then $(f_{N})^{*}(L)\simeq L^{\otimes d(f)^{N}}$ and $f_{N}\circ f_{N^{\prime}}=f_{N}\circ f_{N^{\prime}}=f_{N+N^{\prime}}$. Note that $d_{N}=d(f)^{N}$ and $\{f_{N}\}_{N=1}^{\infty}$ is additive. In this case, $h_{F}(x)=0$ if and only if $x$ is preperiodic with respect to $f$. It seems that Theorem 1.2 is slightly related to the problem posed in [1]. We assume that $f(Y)\subseteq Y$. Then we have the tower
$Y(K)\subseteq Y_{1}(K)\subseteq\cdots\subseteq Y_{N}(K)\subseteq\cdots$
If $Y(K)$ is finite, by (1) in Theorem 1.2, there is $N_{2}$ such that
$Y_{N}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}$
for all $N\geqslant N_{2}$, so the tower is stabilized because $\{x\in X(K)\mid h_{F}(x)=0\}$ is finite.
Remark 4.7.
The above Example 4.2, Example 4.3, Example 4.4 and Example 4.5 are exceptional from the dynamics point of view, but Example 4.6 is the general iterative setting. In this sense, the actual analogue of Fermat’s conjecture is very restricted in the dynamical systems.
Remark 4.8.
In Example 4.2, we assume that $K={\mathbb{Q}}$, $n=2$ and $Y$ is defined by $X_{0}+X_{1}-X_{2}=0$. Note that $Y_{N}$ is given by $X_{0}^{N}+X_{1}^{N}-X_{2}^{N}$. If $Y_{N}$ has Fermat’s property, then, for $(x_{0}:x_{1}:x_{n})\in Y_{N}(K)$, we can find $t\in{\mathbb{Q}}^{\times}$ such that $tx_{0},tx_{1},tx_{2}\in\{0,1,-1\}$. Thus we can easily see that $x_{0}x_{1}x_{2}=0$, and hence Fermat’s conjecture for $N$ follows.
Remark 4.9.
If $\dim Y=1$, then $\dim Y_{N}=1$ because $f_{N}$ is finite. Thus if there is $N_{0}$ such that every irreducible component of $Y_{N}$ has the genus $\geqslant 2$ for any integer $N\geqslant N_{0}$, then, by Faltings theorem (c.f. [6]), $Y_{N}(K)$ is finite for $N\geqslant N_{0}$.
Remark 4.10.
If $\{f_{N}\}_{N=N_{0}}^{\infty}$ is a multiplicative system of endomorphisms polarized by $L$, then, for a fixed integer $m\geqslant 2$, $\{g_{N}\}_{N=N_{1}}^{\infty}$ given by $g_{N}=f_{m^{N}}$ is additive, where $N_{1}$ is an integer with $m^{N_{1}}\geqslant N_{0}$.
Remark 4.11.
Theorem 1.2 was previously discussed in the following cases:
  1. (1)
    $K={\mathbb{Q}}$, $X={\mathbb{P}}^{2}$, $f_{N}(x_{0}:x_{1}:x_{2})=(x_{0}^{N}:x_{1}^{N}:x_{2}^{N})$ and $Y$ is given by $X_{0}+X_{1}-X_{2}$ (M. Filaseta [7], A. Granville [8], D. R. Heath-Brown [9]).
  2. (2)
    $K$ is a number field, $X={\mathbb{P}}^{2}$, $f_{N}(x_{0}:x_{1}:x_{2})=(x_{0}^{N}:x_{1}^{N}:x_{2}^{N})$ and $Y$ is given by $X_{0}+X_{1}-X_{2}$ (H. Ikoma, S. Kawaguchi and A. Moriwaki [10]).
  3. (3)
    $K$ is an arithmetic function field, $X={\mathbb{P}}^{2}$, $f_{N}(x_{0}:x_{1}:x_{2})=(x_{0}^{N}:x_{1}^{N}:x_{2}^{N})$ and $Y$ is a strongly non-degenerated curve, that is, $Y_{N}$ is smooth over $K$ for all NN (H. Chen and A. Moriwaki [4]).
Proof of Theorem 1.2.
Let $S$ be a finite subset of $X(K)$. As $\lim_{m\to\infty}d_{m}=\infty$ by (1) of Proposition 3.1, there is $m_{0}$ such that
$d_{m}a(h_{F})>\max\{h_{F}(y)\mid y\in S\}$
for all $m\in{\mathbb{Z}}_{\geqslant m_{0}}$.
(1) We set $S=Y(K)$. Then, for $N\in{\mathbb{Z}}_{\geqslant m_{0}}$, as $f_{N}(Y_{N}(K))\subseteq S$ and $f_{N}\circ f_{N_{0}}=f_{N_{0}}\circ f_{N}$, by Lemma 5.2, $Y_{N}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}$.
(2) Note that $f^{-1}_{N}(Y_{N_{1}})=Y_{N+N_{1}}$. Thus the assertion follows from (1).
(3) For any prime number $p\geqslant p_{0}$, we set $S=Y_{p}(K)$. If $x\in Y_{mp}(K)$, then one has $f_{m}(x)\in Y_{p}(K)$ because $f_{p}(f_{m}(x))=f_{mp}(x)\in Y(K)$. Moreover, $f_{m}\circ f_{N_{0}}=f_{N_{0}}\circ f_{m}$. Thus, by Lemma 5.2, $Y_{mp}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}$ for $m\in{\mathbb{Z}}_{\geqslant m_{0}}$. Therefore, by [10, Lemma 5.16] or Lemma 5.6, we have the assertion. ∎
5. Product case (multi-indexed version)
Let $X_{1},\ldots,X_{n}$ be geometrically integral projective schemes over $K$ and $L_{1},\ldots,L_{n}$ be ample line bundles on $X_{1},\ldots,X_{n}$, respectively. Let $f_{1},\ldots,f_{n}$ be endomorphisms of $X_{1},\ldots,X_{n}$, respectively. We assume that, for each $i\in\{1,\ldots,n\}$, $f_{i}^{*}(L_{i})\simeq L_{i}^{\otimes d_{i}}$ for some integer $d_{i}\geqslant 2$. Let $\varphi_{i}$ be the global compatification of $f_{i}^{*}(L_{i})\simeq L_{i}^{d_{i}}$ for each $i$. Let $X=X_{1}\times\cdots\times X_{n}$, $L=p_{1}^{*}(L_{1})\otimes\cdots\otimes p_{n}^{*}(L_{n})$ and $f=f_{1}\times\cdots\times f_{n}$, that is, $f(x_{1},\ldots,x_{n})=(f_{1}(x_{1}),\ldots,f_{n}(x_{n}))$ for $(x_{1},\ldots,x_{n})\in X$. Moreover, let $p_{i}:X\to X_{i}$ be the projection to the $i$-th factor. Note that $f^{*}(L)\simeq p_{1}^{*}(L_{1})^{\otimes d_{1}}\otimes\cdots\otimes p_{n}^{*}(L_{n})^{\otimes d_{n}}$, and $\varphi=p_{1}^{*}(\varphi_{1})\otimes\cdots\otimes p_{n}^{*}(\varphi_{n})$ yields the global compatification of $f^{*}(L)\simeq p_{1}^{*}(L_{1})^{\otimes d_{1}}\otimes\cdots\otimes p_{n}^{*}(L_{n})^{\otimes d_{n}}$, that is, we have the isometry
$f_{\omega}^{*}(L_{\omega},\varphi_{\omega})\simeq p_{1,\omega}^{*}(L_{1,\omega},\varphi_{1,\omega})^{\otimes d_{1}}\otimes\cdots\otimes p_{n,\omega}^{*}(L_{n,\omega},\varphi_{n,\omega})^{\otimes d_{n}}$
for each $\omega\in\Omega$.
Proposition 5.1.
For $x_{1}\in X_{1}(K),\ldots,x_{n}\in X_{n}(K)$,
$\begin{cases}h_{f}(x_{1},\ldots,x_{n})=h_{f_{1}}(x_{1})+\cdots+h_{f_{n}}(x_{n}),\\ h_{f}(f(x_{1},\ldots,x_{n}))=d_{1}h_{f_{1}}(x_{1})+\cdots+d_{n}h_{f_{n}}(x_{n}).\end{cases}$
Proof.
Indeed,
$h_{f}(x_{1},\ldots,x_{n})$ $=h_{(p_{1}^{*}(L_{1})\otimes\cdots\otimes p_{n}^{*}(L_{n}),p_{1}^{*}(\varphi_{1})\otimes\cdots\otimes p_{n}^{*}(\varphi_{n}))}(x_{1},\ldots,x_{n})$
$=h_{(p_{1}^{*}(L_{1}),p_{1}^{*}(\varphi_{1}))}(x_{1},\ldots,x_{n})+\cdots+h_{(p_{n}^{*}(L_{n}),p_{n}^{*}(\varphi_{n}))}(x_{1},\ldots,x_{n})$
$=h_{f_{1}}(x_{1})+\cdots+h_{f_{n}}(x_{n}).$
Moreover,
$h_{f}(f(x_{1},\ldots,x_{n}))$ $=h_{f}(f_{1}(x_{1}),\ldots,f_{n}(x_{n}))=h_{f_{1}}(f_{1}(x_{1}))+\cdots+h_{f_{n}}(f_{n}(x_{n}))$
$=d_{1}h_{f_{1}}(x_{1})+\cdots+d_{n}h_{f_{n}}(x_{n}),$
as required. ∎
Lemma 5.2.
Let $S$ be a finite subset of $X(K)$. For each $i\in\{1,\ldots,n\}$, let $g_{i}:X_{i}\to X_{i}$ be an endomorphism of $X_{i}$ such that $g_{i}$ are polarized by $L_{i}$, $d(g_{i})\in{\mathbb{Z}}_{\geqslant 2}$, and $f_{i}\circ g_{i}=g_{i}\circ f_{i}$. We set $g=g_{1}\times\cdots\times g_{n}$. If $\min\{d(g_{i})a(h_{f_{i}})\mid i\in\{1,\ldots,n\}\}>\max\{h_{f}(s)\mid s\in S\}$ and $g(x)\in S$ for $x\in X(K)$, then $h_{f}(x)=0$.
Proof.
We set $x=(x_{1},\ldots,x_{n})$ and $I=\{i\in\{1,\ldots,n\}\mid h_{f_{i}}(x_{i})>0\}$. If we assume $I\not$$\emptyset$, then
$\max\{h_{f}(x)\mid x\in S\}$ $\geqslant h_{f}(g(x))=\sum_{i=1}^{n}h_{f_{i}}(g_{i}(x_{i}))=\sum_{i=1}^{n}h_{g_{i}}(g_{i}(x_{i}))$
$=\sum_{i=1}^{n}d(g_{i})h_{g_{i}}(x_{i})=\sum_{i=1}^{n}d(g_{i})h_{f_{i}}(x_{i})$
$=\sum_{i\in I}d(g_{i})h_{f_{i}}(x_{i})\geqslant\sum_{i\in I}d(g_{i})a(h_{f_{i}})$
$\geqslant\min\{d(g_{i})a(h_{f_{i}})\mid i\in\{1,\ldots,n\}\},$
which is a contradiction. ∎
For $I=(i_{1},\ldots,i_{n})\in({\mathbb{Z}}_{\geqslant 1})^{n}$ and $1\leqslant k\leqslant n$, we denote $i_{k}$ by $I(k)$. We define $\min I$ to be $\min\{I(k)\mid k\in\{1,\ldots,n\}\}$. For $I,J\in({\mathbb{Z}}_{\geqslant 1})^{n}$, we define $I\leqslant J$, $I+J$ and $I\cdot J$ to be
$\begin{cases}I\leqslant J\ \Longleftrightarrow\ I(1)\leqslant J(1),\ldots,I(n)\leqslant J(n),\\ I+J=(I(1)+J(1),\ldots,I(n)+J(n)),\\ I\cdot J=(I(1)\cdot J(1),\ldots,I(n)\cdot J(n)).\end{cases}$
Moreover, we say $I$ is prime if $I(k)$ is a prime number for every $k\in\{1,\ldots,n\}$.
Let $F_{1}=\{f_{1,N}\}_{N=N_{1}}^{\infty},\ldots,F_{n}=\{f_{n,N}\}_{N=N_{n}}^{\infty}$ be systems of endomorphisms on $X_{1},\ldots,X_{n}$, respectively such that $f_{k,N}$ is polarized by $L_{k}$, that is, $f^{*}_{k,N}(L_{k})\simeq L^{\otimes d_{k,N}}$ ($d_{k,N}\in{\mathbb{Z}}_{\geqslant 2}$). We set $I_{0}=(N_{1},\ldots,N_{n})$. For $I\geqslant I_{0}$, the endomorphism $f_{1,I(1)}\times\cdots\times f_{n,I(n)}$ on $X$ is denoted by $f_{I}$, and the collection $\{f_{I}\}_{I\in({\mathbb{Z}}_{\geqslant 1})^{n}_{\geqslant N_{0}}}$ is denoted by $F$. Note that the following properties are easily verified:
  1. (1)
    For $I,I^{\prime}\in({\mathbb{Z}}_{\geqslant 1})^{n}_{\geqslant N_{0}}$, $f_{I}\circ f_{I^{\prime}}=f_{I^{\prime}}\circ f_{I}$.
  2. (2)
    $f_{I}^{*}(p_{1}^{*}(L_{1})\otimes\cdots\otimes p_{n}^{*}(L_{n}))=p_{1}^{*}(L_{1})^{\otimes d_{1,I(1)}}\otimes\cdots\otimes p_{n}^{*}(L_{n})^{\otimes d_{n,I(n)}}$.
Moreover, note that the height function $h_{f_{I}}$ does not depend on the choice of $I$, so it is denoted by $h_{F}$. Let $Y$ be an equidimensional subscheme of $X$, and $Y_{I}=f^{-1}_{I}(Y)$. We say $Y_{I}$ has Fermat’s property over $K$ if
$Y_{I}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}.$
Corollary 5.3.
If $F$ is multiplicative, that is, $f_{I}\circ f_{I^{\prime}}=f_{I^{\prime}}\circ f_{I}=f_{I\cdot I^{\prime}}$ for all $I,I^{\prime}\in({\mathbb{Z}}_{\geqslant 1})^{n}_{\geqslant N_{0}}$ and $Y_{I}(K)$ is finite, then there is $M_{0}$ such that $Y_{I\cdot M}$ has Fermat’s property for all $M\geqslant M_{0}$.
Proof.
This is a consequence of Lemma 5.2. ∎
Conjecture 5.4 (Multi-indexed version of generalized Fermat’s conjecture).
If there is $I_{1}\geqslant I_{0}$ such that $Y_{I}(K)$ is finite for any $I\geqslant I_{1}$, then does there exist $I_{2}\geqslant I_{0}$ such that $Y_{I}$ has Fermat’s property over $K$ for any $I\geqslant I_{2}$?
One has the following as an evidence of the above conjecture.
Theorem 5.5.
If $F$ is multiplicative and there is $I_{1}\geqslant I_{0}$ such that $Y_{I}(K)$ is finite for all $I\geqslant I_{1}$, then
$\lim_{m_{1}\to\infty,\ldots,m_{n}\to\infty}\frac{\#\left\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\ \left|\ \begin{array}[]{l}\text{$I_{0}\leqslant I\leqslant(m_{1},\ldots,m_{n})$ and }\\ \text{$Y_{I}$ has Fermat's property over $K$}\end{array}\right.\right\}}{m_{1}\cdots m_{n}}=1.$
Proof.
This is a consequence of the above corollary and the following lemma. ∎
Lemma 5.6.
Let $\Sigma$ be a subset of $({\mathbb{Z}}_{\geqslant 1})^{n}$. We assume that there is a positive integer $p_{0}$ with the following properties: for any prime $P\in({\mathbb{Z}}_{\geqslant 1})^{n}$ with $\min P\geqslant p_{0}$ (i.e., $P(i)$ is a prime number and $P(i)\geqslant p_{0}$ for every $i$), there is $M^{\prime}\in({\mathbb{Z}}_{\geqslant 1})^{n}$ such that $P\cdot M\in\Sigma$ for any $M\geqslant M^{\prime}$. Then
$\lim_{m_{1}\to\infty,\ldots,m_{n}\to\infty}\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$I\in\Sigma$ and $I\leqslant(m_{1},\ldots,m_{n})$}\}}{m_{1}\cdots m_{n}}=1.$
Proof.
Since the Riemann zeta function has a pole at $1$, for a positive number $\varepsilon$ with $0<\varepsilon\leqslant 1$, let $p_{1},\ldots,p_{\ell}$ be prime numbers such that $p_{0}\leqslant p_{1}<\cdots<p_{\ell}$ and $(1-1/p_{1})\cdots(1-1/p_{\ell})\leqslant\varepsilon/(4n)$. For each $Q\in\{p_{1},\ldots,p_{\ell}\}^{n}$, by our assumption, there exists $M_{Q}\in({\mathbb{Z}}_{\geqslant 1})^{n}$ such that $Q\cdot M\in\Sigma$ for all $M\geqslant M_{Q}$.
Claim 5.6.1.
If we set $e=p_{1}\cdots p_{\ell}$, then we can find $M_{0}\in{\mathbb{Z}}_{\geqslant 1}^{n}$ such that
$\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$I\geqslant M_{0}$ and $\operatorname{GCD}(I(i),e)\not=1$ for all $i\in\{1,\ldots,n\}$}\}\subseteq\Sigma$
Proof.
We set $M_{0}=\sum_{Q\in\{p_{1},\ldots,p_{\ell}\}^{n}}QM_{Q}$. Let $I$ be an element of $({\mathbb{Z}}_{\geqslant 1})^{n}$ such that $I\geqslant M_{0}$ and $\operatorname{GCD}(I(i),e)\not=1$ for all $i\in\{1,\ldots,n\}$. Then, for each $i\in\{1,\ldots,n\}$, there is $q_{i}\in\{p_{1},\ldots,p_{\ell}\}$ such that $q_{i}\mid I(i)$. Thus, if we set $Q=(q_{1},\ldots,q_{n})$, then $I=Q\cdot M$ for some $M\in({\mathbb{Z}}_{\geqslant 1})^{n}$. Moreover,
$QM_{Q}\leqslant M_{0}\leqslant I=Q\cdot M,$
so $M\geqslant M_{Q}$, and hence $I\in\Sigma$. ∎
We set $M_{0}=(a_{1},\ldots,a_{n})$. Then we have the following:
Claim 5.6.2.
If $m_{1}\geqslant a_{1},\ldots,m_{n}\geqslant a_{n}$, then
$\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\setminus\Sigma\mid I\leqslant(m_{1},\ldots,m_{n})\}}{m_{1}\cdots m_{n}}\leqslant\sum_{i=1}^{n}\frac{a_{i}-1}{m_{i}}+(1-1/p_{1})\cdots(1-1/p_{\ell})\sum_{i=1}^{n}\frac{m_{i}+e}{m_{i}}$
Proof.
By the above claim,
$({\mathbb{Z}}_{\geqslant 1})^{n}\setminus\Sigma$ $\subseteq\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid I\not\geqslant M_{0}\}\cup\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$\exists i\in\{1,\ldots,n\}$ $\operatorname{GCD}(I(i),e)=1$}\}$
$\kern-20.00003pt\subseteq\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$\exists i\in\{1,\ldots,n\}$ $I(i)<a_{i}$}\}\cup\bigcup_{i=1}^{n}\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$\operatorname{GCD}(I(i),e)=1$}\}$
$\kern-20.00003pt\subseteq\bigcup_{i=1}^{n}\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$I(i)<a_{i}$}\}\cup\bigcup_{i=1}^{n}\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$\operatorname{GCD}(I(i),e)=1$}\}$
Thus, by [10, the proof of Lemma 5.16], if $m_{1}\geqslant a_{1},\ldots,m_{n}\geqslant a_{n}$, then
$\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\setminus\Sigma\mid I\leqslant(m_{1},\ldots,m_{n})\}}{m_{1}\cdots m_{n}}\\ \kern-100.00015pt\leqslant\sum_{i=1}^{n}\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$I(i)<a_{i}$ and $I\leqslant(m_{1},\ldots,m_{n})$}\}}{m_{1}\cdots m_{n}}\\ \kern 50.00008pt+\sum_{i=1}^{n}\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$\operatorname{GCD}(I(i),e)=1$ and $I\leqslant(m_{1},\ldots,m_{n})$}\}}{m_{1}\cdots m_{n}}\\ \leqslant\sum_{i=1}^{n}\frac{a_{i}-1}{m_{i}}+\sum_{i=1}^{n}\frac{(1-1/p_{1})\cdots(1-1/p_{\ell})(m_{i}+e)}{m_{i}},$
as required. ∎
We choose positive integers $b_{1},\ldots,b_{n}$ such that $(a_{i}-1)/b_{i}\leqslant\varepsilon/(2n)$ and $e/b_{i}\leqslant\varepsilon$. Then, by the above claim, if $m_{i}\geqslant\max\{a_{i},b_{i}\}$ for all $i\in\{1,\ldots,n\}$, we have
$\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\setminus\Sigma\mid I\leqslant(m_{1},\ldots,m_{n})\}}{m_{1}\cdots m_{n}}\leqslant\varepsilon/2+(\varepsilon/4)(1+\varepsilon)\leqslant\varepsilon,$
because $0<\varepsilon\leqslant 1$, and hence the assertion follows. ∎
Example 5.7.
Let $f_{(A,A^{\prime})}$ $(A,A^{\prime}\geqslant 2$) be the endomorphism on ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}(=\operatorname{Proj}(K[X_{0},X_{1}])\times\operatorname{Proj}(K[Y_{0},Y_{1}]))$ defined to be
$f_{(A,A^{\prime})}((x_{0}:x_{1})\times(y_{0}:y_{1}))=(x_{0}^{A}:x_{1}^{A})\times(y_{0}^{A^{\prime}}:y_{1}^{A^{\prime}}).$
Let $p_{i}:{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}\to{\mathbb{P}}^{1}$ be the projection to the $i$-th factor, and $L=p_{1}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))\otimes p_{2}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))$. Note that $f_{(A,A^{\prime})}\circ f_{(B,B^{\prime})}=f_{(B,B^{\prime})}\circ f_{(A,A^{\prime})}=f_{(AB,A^{\prime}B^{\prime})}$ (multiplicative). Let $Y$ be a subvariety of ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ given by
$\alpha X_{0}Y_{0}+\beta X_{1}Y_{1}+\gamma X_{0}Y_{1}+\delta X_{1}Y_{0}=(X_{0}\ X_{1})\Pi\begin{pmatrix}Y_{0}\\ Y_{1}\end{pmatrix}=0\quad(\alpha,\beta,\gamma,\delta\in K),$
where $\Pi=\begin{pmatrix}\alpha&\gamma\\ \delta&\beta\end{pmatrix}$. Then $Y_{(A,A^{\prime})}$ is given by
$\alpha X_{0}^{A}Y_{0}^{A^{\prime}}+\beta X_{1}^{A}Y_{1}^{A^{\prime}}+\gamma X_{0}^{A}Y_{1}^{A^{\prime}}+\delta X_{1}^{A}Y_{0}^{A^{\prime}}=(X^{A}_{0}\ X^{A}_{1})\Pi\begin{pmatrix}Y^{A^{\prime}}_{0}\\ Y^{A^{\prime}}_{1}\end{pmatrix}=0.$
We assume that $\alpha\beta\gamma\delta\not=0$ and $\det\Pi\not=0$. Then it is easy to see that $Y_{(A,A^{\prime})}$ is smooth over $K$ for all $A,A^{\prime}\geqslant 2$, so its genus is $(A-1)(A^{\prime}-1)$. Thus, Theorem 5.5 holds in this case.